OpenStudy (anonymous):

Solve: |3r - 12| = 0

OpenStudy (anonymous):

Absolute values give two answers. First, 3r - 12 = +0 3r = 12 r = 4 Second, 3r - 12 = -0 r = 4 In this case r = 4 because positive and negative zero are identical. However, it is common for the positive and negative versions to be significant.

OpenStudy (anonymous):

\[3r + 12 = \pm0\] 4,4